OpenStudy (loser66):

Suppose that x and y are vectors and M is a subspace in a vector space V Let H be the subspace spanned by M and x Let K be the subspace spanned by M and y. Prove that if y is in H but not in M, then x is in K. Please, help

OpenStudy (loser66):

H = Span (M +x) K = Span ( M + y) case 1: x \(\in M\), y \(\in M\) case 2: x \(\in M\) , y \(\notin M\) case 3: x\(\notin M\) , y \(\in M\) case4: both x , y \(\notin M\) Is there any else? please, help